given
F(x) = 2x -4
G(x) = 5 + x^2
find (F o G)(x)

Answer by venugopalramana(3286)

given
F(x) = 2x -4
G(x) = 5 + x^2=Y SAY
find (F o G)(x)=F{G(X)}=F(Y)=2Y-4=2(5+X^2-4)=2X^2+2